how to solve this equation by using the square root porperty. (1-x)^2 = 4
Answer by Cromlix(4381)

(1 - x)^2 = 4
Square root both sides
1 - x = sqrt4
1 - x = 2
- x = 2 - 1
- x = 1
x = -1
Proof
(1 - x)^2 = 4
(1 - (-1)^2 = 4
(1 + 1)^2 = 4
(2)^2 = 4
4 = 4
Hope this helps.
:-)